Consider the following reaction.
N2(g) + O2(g) 2 NO(g)
If the equilibrium partial pressures of N2, O2, and NO are 0.15 atm, 0.33 atm, and 0.050 atm, respectively, at 2200°C, what is KP?
Answer #1
Concepts and reason

The concept used to solve this problem is determined the Equilibrium constant.

Fundamentals

At equilibrium, the rate of forward reaction and the rate of reverse reaction; both are same, and will not change with time, than the value of the reaction quotient is called equilibrium constant.

The equilibrium constant indicate by the sign of

At equilibrium, the value of the reaction quotient is called equilibrium constant.

is a ratio of the equilibrium concentrations of the product to the equilibrium concentration of the reactants.

For example the equilibrium constant is calculated as follows:

aA + bB
C + DD

The equilibrium constant ()

Keq=[Products][Reactants]=[C]c[D]d[A]a[B]b\begin{array}{l}\\{K_{{\rm{eq}}}} = \frac{{{\rm{[Products]}}}}{{[{\rm{Reactants}}]}}\\\\{\rm{ }} = {\rm{ }}\frac{{{{\left[ {{\rm{ C}}} \right]}^c}{{\left[ {{\rm{ D}}} \right]}^d}}}{{{{\left[ {\rm{A}} \right]}^a}{{\left[ {\rm{B}} \right]}^b}}}\\\end{array}

In the above reaction; a, b, c and d are coefficients

The equilibrium constant can be expressed in terms of partial pressures of the reactants and products.

According to the ideal gas law:

PV = nRT
P=Conc.x RT
Poc Conc.

Here, is the pressure, is the volume, is the gas constant, is the temperature and is the number of moles. is the concentration.

Therefore; the equilibrium constant; for gaseous reaction;

PEPP
PAPER

The given balance reaction is as follows:

N, (g) + O2(g) = 2 NO (8)

The equilibrium constant; for gaseous reaction;

K = =

According to the problem the equilibrium partial pressures of = 0.15 atm, partial pressures of = 0.33 atm, and partial pressures of = 0.050 atm, at 2200°C

The equilibrium constant () is calculated as follows:

Kp=PNO2PN21PO21=[0.050]2[0.15]1[0.33]1=0.00250.0495=0.0505\begin{array}{l}\\{\rm{ }}{K_{\rm{p}}} = {\rm{ }}\frac{{{P_{{\rm{NO}}}}^2}}{{{P_{{{\rm{N}}_{\rm{2}}}}}^1\;{P_{{{\rm{O}}_{\rm{2}}}}}^1}}\\\\{\rm{ }} = {\rm{ }}\frac{{{{\left[ {{\rm{0}}{\rm{.050}}} \right]}^2}}}{{{{\left[ {{\rm{0}}{\rm{.15}}} \right]}^1}{{\left[ {{\rm{0}}{\rm{.33}}} \right]}^1}}}\\\\{\rm{ }} = {\rm{ }}\frac{{{\rm{0}}{\rm{.0025}}}}{{0.0495}}\\\\{\rm{ }} = {\rm{ 0}}{\rm{.0505}}\\\end{array}

Ans:

Thus, for given reaction is 0.0505
.